OpenStudy (anonymous):

600-kg car is going over a curve with a radius of 120 meters that is banked at an angle of 25 degrees with a speed of 30 meters per second. The coefficient of static friction between the car and the road is 0.3. What is the normal force exerted by the road on the car?

OpenStudy (anonymous):

Normal reaction of road is centrifugal force by sinus of banked angle static frict coef not to be taken into acc as normal force to be calculated; \[m.r.w ^{2}.sin25 where w is angular speed.\] angular speed\[30/120=v/r\]=0.25 rd/s\[600\times120\times0,25^{2}\times \sin 25\] force:1900N or 193.7kgf